OpenStudy (anonymous):

Graph the line. 3x - 6y = 18

jimthompson5910 (jim_thompson5910):

when you plug in x = 0, what do you get when you solve for y?

OpenStudy (anonymous):

(-3x) 3x - 6y = 18 (- 3x) -6y = 18 - 3x

OpenStudy (anonymous):

Is that what I do?

jimthompson5910 (jim_thompson5910):

3x - 6y = 18 3*0 - 6y = 18 ... replace x with 0 (aka, plug in x = 0) 0 - 6y = 18 -6y = 18 y = ???

OpenStudy (anonymous):

Oh, -3

jimthompson5910 (jim_thompson5910):

yep, so (0,-3) is on the line

jimthompson5910 (jim_thompson5910):

when y = 0, what is x?

OpenStudy (anonymous):

3x - 0y = 18 3x = 18 /3 /3 x = 6

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

good, x = 6, y = 0 so the point (6,0) is on the line

OpenStudy (anonymous):

So 6,3?

jimthompson5910 (jim_thompson5910):

so we know these two points lie on the line y-intercept (0,-3) x-intercept (6,0)

jimthompson5910 (jim_thompson5910):

plot those two points, then draw a line through them

OpenStudy (anonymous):

Oh, thanks,

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

c: